Please help me solve this equation: The question says,Solve the following equations for two unknowns (simultaneously)
a) 2p+5q=6
q-p=4
B) -3(g+2c)=-27
8g-12c=44

Answer by Alan3354(69443)

a) 2p+5q=6
q-p=4 --> q = p+4
Since the 2nd eqn is in terms of q, use substition
2p+5q=6
2p + 5(p+4) = 6
7p+20 = 6
7p=-14
p = -2
---------
q = +2
-----------
B) -3(g+2c)=-27 --> g+2c = 9 --> 2g+4c = 18
8g-12c=44 --> 2g-3c = 11
--------------
2g+4c = 18
2g-3c = 11
------------ Subtract
0g+7c = 7
c = 1
------
g = 7